# 1in1in1in1in

### v8

@@ 1 Introduction
As AI applications demand greater compute power, efficiency may be improved via better chip design. The Nature paper [1] by Google researchers, published in June 2021, was advertised as a chip-design breakthrough using Machine Learning (ML). It addressed a challenging problem to optimize locations of circuit components on a chip and described applications to five Tensor Processing Unit (TPU) chip blocks, implying that no better methods were available at the time in academia or industry. The paper broadened the claims beyond chip design to suggest that Reinforcement Learning (RL) extends state of the art in combinatorial optimization.
@@
Given the earlier exodus of ML researchers from Google, the New York Times and Reuters covered this controversy in 2022 [53, 54]. The reporters found that, well before the Nature submission, Google researchers tasked with checking its claims disputed those claims. Two lead authors of [1] complained of persistent allegations of fraud in their work [55]. In 2022, Google fired the ML researcher who made the allegations [53, 54] and denied publication approval for his paper [5] critical of [1]. The dissenter sued Google for wrongful termination under California whistleblower-protection laws: court documents [58], filed under penalty of perjury, show allegations of fraud and scientific misconduct related to work in [1]. The lawsuit moved ahead in Aug ‘23 [65, 67, 73]. Within months of the 2022 media investigations and the lawsuit, the two lead authors of [1] and a senior coauthor left Google [56]. Spring 2023 media coverage noted alleged misrepresentations by Google to potential cloud-services customers [60], questioned reproducibility of results in the Nature paper [63], and covered UCSD research trying to settle the dispute [64]. The 2021 Nature News & Views article introducing [1] in the same issue urged replication of results of [1]. Given the obstacles to and the results of replication [7], the author retracted the article.
In this work, Section 2 reviews background and the chip-design task solved in [1, 3], it also introduces our secondary sources [4, 5, 6, 7]. Section 3 lists initial suspicions about [1]. Section 4 shows that many of them were later confirmed. Section 5 checks if [1] improved the State of the Art (SOTA). Section 6 outlines technical responses by authors of [1] to critiques. Section 7 outlines how the work from [1] may be used in practice. Section 8 draws conclusions about [1] and notes policy implications.
@@
4 proprietary TPU blocks 20 proprietary TPU blocks
@@
2×2$\times$2 × MemPool (public) [6, 7]
@@ 2 Background
Circuit placement as an optimization task. After $(x,y)$ locations of all components are known, wires that connect components’ I/O pins are routed. Routes impact chip metrics (for power, timing/speed, etc). The optimization of $(x,y)$ locations starts with simplified estimates of wirelength without wire routes. Pin locations $(x_{1},y_{1})$ and $(x_{2},y_{2})$ may be connected by horizontal and vertical wire segments in many ways, but the shortest route length is $|x_{1}-x_{2}|+|y_{1}-y_{2}|$. For multiple locations $\{(x_{i},y_{i})\}_{i}$, this estimate generalizes to
@@
Widely adopted optimization frameworks for placement do not use ML [29, 30, 39, 51, 48] and can be classified as: ($i$) Simulated Annealing (SA), ($ii$) partitioning-driven, and ($iii$) analytical. SA, developed in the 1980s [10, 12, 13] and dominant through the mid-1990s [14, 15, 16], starts with an initial layout (e.g., random) and alters it by a sequence of actions, such as component moves and swaps, of prescribed length. To improve the final result, some actions may sacrifice quality. SA excels on smaller layouts (up to 100K placeable components) but takes long time for large layouts. Partitioning-driven methods [18, 24, 25, 27, 26] view the circuit connectivity (the netlist) as a hypergraph and use established software packages to subdivide it into partitions with more connections within the partitions (not between). These methods run faster than SA, capture global netlist structures, and were dominant for some 10 years. Yet, the mismatch between partitioning and placement objectives (Equation 1) leaves room for improvement [26]. Analytical methods approximate Equation 1 by closed-form functions amenable to established optimization methods. Force-directed placement [11] from the 1980s models nets by springs and finds component locations to balance out spring forces [51]. In the 2000s, advanced analytical placement techniques attained superiority [29, 39, 44, 48] on all large modern public benchmark sets, including those with macros and routing data  [28, 31, 32, 33, 35, 44]. RePlAce [44] from UCSD runs much faster than SA and partitioning-based methods, but lags in quality on small netlists.
@@
Among secondary sources discussed in the context of [1] (Table 1), we prefer scholarly papers [4, 5, 7] and draw on open-source repositories and FAQs as needed [3, 6]. We crosscheck claims from three nonoverlapping groups of researchers: those associated with [1], the Stronger Baselines paper [5] and the UCSD paper [7]. Consistent claims from different groups are even more trustworthy when backed by numerous benchmarks. Both Google Team 2 and the UCSD team included highly-cited experts on floorplanning and placement with extensive publication records and several key references cited in [1] (such as [32, 39, 44] and others), as well as experience developing academic and commercial floorplanning and placement tools beyond Google.
@@ 3 Initial doubts
While the Nature paper [1] was sophisticated and impressive, its research plan had notable shortfalls. For one, proposed reinforcement learning (RL) was presented as being capable of broader combinatorial optimization (a field that includes puzzle-like tasks such as the Traveling Salesperson Problem, Vertex Cover, Bin Packing). But instead of illustrating this with key problem formulations and easy-to-configure test examples [47], it solved a specialty task (macro placement for chip design) for proprietary Google TPU circuit design blocks, giving results on 5 blocks out of many more available. The RL formulation did not track chip metrics and optimized a simplified proxy function that included HPWL (see Section 2), but was not evaluated for pure HPWL optimization on open circuit examples, as is routine in the literature [18, 20, 23, 24, 25, 26, 30, 29, 35, 39, 44, 48]. New ideas in placement are usually evaluated in research contests on industry chip designs released as public benchmarks [48, Section 6.1], e.g., [29, 36, 37, 38, 39, 41]. but [1] neglected these contest benchmarks.
Some aspects of [1] looked suspicious, as it ($i$) did not substantiate several claims and withheld key aspects of experiments, ($ii$) claimed improvements in noisy metrics that the proposed technique did not optimize, ($iii$) relied on techniques with known handicaps that undermined performance in similar circumstances, and (4) may have misconfigured and underreported its baselines. We spell these out in Sections 3.1-3.4 — confirming even a fraction of specific concerns would put the top-line claims and conclusions of [1] in serious doubt.
Refer to captionRefer to caption
Figure 1: Layouts from [5, Figure 2] with macros in red and standard cells in green, locations produced by RL (left) and RePlAce (right) for the ibm10 benchmark from [22]. Limiting macro locations to a coarse grid (left) leads to spreading of small macros (red squares on a grid) and elongates connecting wires: from 27.5 (right) to 44.1 (left) for ibm10 [5, Table 1]). High area utilization and many macros of different sizes make the ICCAD 2004 benchmarks [22] challenging compared to [1] and [2, page 43].
@@ 3.1 Unsubstantiated claims and
U7. For iterative optimization algorithms that improve results over time, fair comparisons show per testcase: ($i$) better quality metrics with equal runtime, ($ii$) better runtime with equal quality or ($iii$) wins for both. [1] offered no such evidence.
@@ 4 Additional evidence
Months after the Nature publication [1], more data became available in [2, 3, 4, 5], followed by the first wave of controversial media coverage [53, 54, 55]. Nature editors released the peer review file for [1] including authors’ rebuttals: in the lengthy back-and-forth with reviewers [2] the authors assured reviewers that macro locations were not modified after placement by RL, confirming coarse-grid placement of macros. Among several contributions, [5] implemented the request of Nature Reviewer #3 [2] and benchmarked Google’s technique on 17 public chip-design examples [22]: prior methods decisively outperformed Google RL. Professors Patrick Madden (SUNY Binghamton) and Jens Lienig (TU Dresden) publicly expressed doubts about the Nature paper [53, 54]. As researchers noted gaps in the Google open-source release [3], such as the grouping (clustering) flow, Google engineers released more code (but not all), prompting more questions (see below). Another year passed, and [6, 7] expanded on the initial suspicions in several ways. They demonstrated that not limiting macro placement allows human designers and commercial EDA tools (separately) to outperform results produced by Google code [3]. [7, Table 2] estimated rank correlation of the proxy cost function optimized by RL to chip metrics used in [1, Table 1], and [7, Table 3] estimated the mean and standard deviation for chip metrics after RL-based optimization. We give a summary in Table 2, where rank correlations are low for all chip metrics, while TNS and WNS are noisy. Hence, the optimization of TNS and WNS in [1] relied on a flawed proxy and produced results in [1, Table 1] of dubious statistical significance. We note that $\sigma/|\mu|>0.5$ for TNS on Ariane-NG45 (also on BlackParrot-NG45 in [7, Table 3]). In the second round of critical media coverage, [1] was questioned by Profs. William Swartz (UT Dallas), Patrick Madden (SUNY Binghamton), and Moshe Vardi (Rice) [60, 63].
@@
Rank correlation to RL proxy cost 0.00 0.28 0.05 0.2 0.05
@@
Table 2: Evaluating the soundness of the proxy cost used with RL in [1] and the noisiness of reported chip metrics after RL-based optimization. We summarize data from [7, Table 2] on the Kendall rank correlation of chip metrics to the RL proxy cost and from [7, Tables 3 and 4] on statistics for chip metrics (only Ariane-NG45 design dats are shown, but data for BlackParrot-NG45 shows similar trends). As expected, purely-additive metrics (standard-cell area, routed wirelength and chip power) exhibit low variance, but the TNS and WNS metrics, that measure timing-constraint violations, have high variance.
@@ 4.1 Methods
Undisclosed use of $(x,y)$ locations from commercial tools. Strong evidence and confirmation by Google engineers are mentioned in the UCSD paper [7] that [1] withheld a critical detail. When clustering the input netlist, CT merge in [3]) read in a placement to restructure clusters based on locations. To produce $(x,y)$ locations of macros, [1] used initial $(x,y)$ locations of all circuit components (including macros!) produced by commercial EDA tools from Synopsys [3, Issue #25]. The lead authors of [1] confirmed using this step, claiming it was unimportant [59]. But it improved key metrics by 7-10% in [7]. So, the results in [1] needed algorithmic steps absent in [1], e.g., $(x,y)$ data from commercial software.
@@
Data leakage between training and test data? Per [1], “as we expose the policy network to a greater variety of chip designs, it becomes less prone to overfitting.” But Google Team 1 showed later in [4, Figure 7] that pre-training on “diverse TPU blocks” did not improve quality of results. Pre-training on “previous netlist versions” improved quality somewhat. Pre-training RL and evaluating it on similar designs could be a serious flaw in methodology of [1]. [^6] Such a methodology could help chip designers iterate on design changes faster, but that was not described in [1]. As Google did not release TPU designs or per-design statistics, we cannot compare training and test data.
@@ 4.3 Likely imitations
[1] did not disclose major limitations of its methods but promised success in broader combinatorial optimization. The Ariane design image in [1, Extended Data Figure 4] shows macro blocks of identical sizes: a potential limitation. Yet, [1] does not report basic statistics per TPU block: the number of macros and their shapes, design area utilization, and the fraction of area taken by macros. Based on [2, page 43] and the guidance from Google engineers to the authors of [7], it appears TPU blocks had area utilization on the order of 60%. Commercial chips are often denser, and can use macros of different sizes. Poor performance of Google RL on challenging public benchmarks from [22] used in [5, 7] (illustrated in Figure 1) may indicate undisclosed limitations. Another possible limitation is poor handling of preplaced (fixed) macros, common in industry layouts, but not discussed in [1]. By interfering with preplaced macros, gridding (see H4) can impact usability in practice. Poor performance on public benchmarks may also be due to overfitting to proprietary TPU designs.
$\downarrow$ Designs / Tools $\rightarrow$ Google CT Cadence CMP UCSD SA
@@ 5 Did [1] improve SOTA?
Empirical validation of research ideas typically seeks to improve State of the Art (SOTA) results. Claims of improving TPU designs [1] would be nonsensical if commercial EDA tools for chip placement were known to considerably outperform the proposed methods. [1] claimed that humans produced better results than commercial EDA tools, but gave no substantiation. When license terms complicate publishing comparisons to commercial EDA tools, [^7] The lawsuit [58, 65] alleges that Google obtained better results with commercial tools before Nature submission. one compares to academic software and to other prior methods, with the proviso that small improvements are not compelling.
@@ 5.1 Reproduction attempts
  • Google Team 2 had access to TPU design blocks and evaluated the impact of pre-training in [5]. The impact was small at best. [^8] A consistent conclusion was reported in [4, Figure 7] by Google Team 1 — training on diverse designs does not help, and improvements are seen only when training in earlier versions of the same design. In December 2022 Dr. Jeff Dean, a then Google SVP and the most senior author of [1], also confirmed that RL did well without pre-training [57].
@@
As explained in [5], the implementation of Simulated Annealing used in [1] was handicapped. Removing the handicaps (in the same source code base) improved results. When properly implemented, SA produces better solutions than Google CT/RL [3] using less runtime, when both are given the same proxy cost function. This is shown consistently in [5, 7] on 17 widely used ICCAD 2004 benchmarks [22] and in [7] on several modern design benchmarks. Compared to Google CT/RL [3], SA consistently improves wirelength and power metrics. For circuit timing metrics TNS and WNS, SA produces less noisy results but comparable to RL’s results [7]. Recall that the proxy function optimized by SA and RL does not include timing metrics [1], making any claims improvement in these metrics due to SA or RL dubious.

### v9

@@ 1 Introduction
As AI applications demand greater compute power, efficiency may be improved via better chip design. The Nature paper [1] by Google researchers, published in June 2021, was advertised as a chip-design breakthrough using Machine Learning (ML) [3]. It addressed a challenging problem to optimize locations of circuit components on a chip and described applications to five Tensor Processing Unit (TPU) chip blocks, implying that no better methods were available at the time in academia or industry. The paper broadened the claims beyond chip design to suggest that Reinforcement Learning (RL) extends state of the art in combinatorial optimization.
@@
Given the earlier exodus of ML researchers from Google, the New York Times and Reuters covered this controversy in 2022 [54, 55]. The reporters found that, well before the Nature submission, Google researchers tasked with checking its claims disputed those claims. Two lead authors of [1] complained of persistent allegations of fraud in their research [56]. In 2022 Google fired an internal whistleblower [54, 55] and denied publication approval for a paper [6] written by Google researchers critical of [1]. The whistleblower sued Google for wrongful termination under California whistleblower-protection laws: court documents [59], filed under penalty of perjury, detail allegations of fraud and scientific misconduct related to research in [1]. The lawsuit moved ahead in Aug ‘23 [66, 68, 74]. Within months of the 2022 media investigations and the lawsuit, the two lead authors of [1] and a senior coauthor left Google [57]. Spring 2023 media coverage noted alleged misrepresentations by Google to potential cloud-services customers [61], questioned reproducibility of results in the Nature paper [64], and covered UCSD research trying to settle the dispute [65]. The 2021 Nature News & Views article introducing [1] in the same issue urged replication of results of [1]. Given the obstacles to and the results of replication [8], the author of the article retracted it.
In this work, Section 2 reviews background and the chip-design task solved in [1, 4]. It also introduces secondary sources used [5, 6, 7, 8]. Section 3 lists initial suspicions about [1]. Section 4 shows that many of them were later confirmed. Section 5 checks if [1] improved the State of the Art (SOTA). Section 6 outlines technical responses by authors of [1] to critiques. Section 7 outlines how the work from [1] may be used in practice. Section 8 draws conclusions about [1] and notes policy implications.
@@
4 proprietary TPU blocks 20 proprietary TPU blocks All with numerous macros:
@@
all with numerous macros all with numerous macros 2×2$\times$2 × MemPool (public) [7, 8]
@@ 2 Background
Circuit placement as an optimization task. After $(x,y)$ locations of all components are known, wires that connect components’ I/O pins are routed. Routes impact chip metrics (for power, timing/speed, etc). The optimization of $(x,y)$ locations starts with simplified estimates of wirelength without wire routes. Pin locations $(x_{1},y_{1})$ and $(x_{2},y_{2})$ may be connected by horizontal and vertical wire segments in many ways, but the shortest route length is $|x_{1}-x_{2}|+|y_{1}-y_{2}|$. For multiple pin locations $\{(x_{i},y_{i})\}_{i}$, this estimate generalizes to
@@
Widely adopted optimizations for placement do not use ML [30, 31, 40, 52, 49] and can be classified as: ($i$) Simulated Annealing (SA), ($ii$) partitioning-driven, and ($iii$) analytical. SA, developed in the 1980s [11, 13, 14] and dominant through the mid-1990s [15, 16, 17], starts with an initial layout (e.g., random) and alters it by a sequence of actions, such as component moves and swaps, of prescribed length. To improve the end result, some actions may sacrifice quality to escape local minima. SA excels on smaller layouts (up to 100K components) but takes a long time for large layouts. Partitioning-driven methods [19, 25, 26, 28, 27] view the circuit connectivity (the netlist) as a hypergraph and use established software packages to subdivide it into partitions with more connections within the partitions (not between). These methods run faster than SA, capture global netlist structures, and were dominant for some 10 years. Yet, the mismatch between partitioning and placement objectives (Equation 1) leaves room for improvement [27]. Analytical methods approximate Equation 1 by closed-form functions amenable to established optimization methods. Force-directed placement [12] from the 1980s models nets by springs and finds component locations to balance out spring forces [52]. In the 2000s, advanced analytical placement techniques attained superiority [30, 40, 45, 49] on all large modern public benchmark sets, including those with macros and routing data  [29, 32, 33, 34, 36, 45]. RePlAce [45] from UCSD is much faster than SA and partitioning methods, but lags in quality on small netlists.
@@
Among secondary sources discussed in the context of [1] (Table 1), we prefer scholarly papers [5, 6, 8] and draw on open-source repositories and FAQs as needed [4, 7]. Here all benchmark sets have hundreds of macros per design, compared to only a handful in sets such as ISPD 2015. We crosscheck claims from three nonoverlapping groups of researchers: those associated with [1], the Stronger Baselines paper [6] and the UCSD paper [8]. Consistent claims from different groups are even more trustworthy when backed by numerous benchmarks. Both Google Team 2 and the UCSD team included highly-cited experts on floorplanning and placement with extensive publication records and several key references cited in [1] (such as [33, 40, 45] and others), as well as experience developing academic and commercial floorplanning and placement tools beyond Google.
@@ 3 Initial doubts
While the Nature paper [1] was sophisticated and impressive, its research plan had notable shortfalls. For one, proposed reinforcement learning (RL) was presented as being capable of broader combinatorial optimization (a field that includes puzzle-like tasks such as the Traveling Salesperson Problem, Vertex Cover, Bin Packing). But instead of illustrating this with key problem formulations and easy-to-configure test examples [48], it solved a specialty task (macro placement for chip design) for proprietary Google TPU circuit design blocks, giving results on 5 blocks out of many more available. The RL formulation did not track chip metrics and optimized a simplified proxy function that included HPWL (see Section 2), but was not evaluated for pure HPWL optimization on open circuit examples, as is routine in the literature [19, 21, 24, 25, 26, 27, 31, 30, 36, 40, 45, 49]. New ideas in placement are usually evaluated in research contests on industry chip designs released as public benchmarks [49, Section 6.1], e.g., [30, 37, 38, 39, 40, 42]. but [1] neglected these contest benchmarks.
Some aspects of [1] looked suspicious, as it ($i$) did not substantiate several claims and withheld key aspects of experiments, ($ii$) claimed improvements in noisy metrics that the proposed technique did not optimize, ($iii$) relied on techniques with known handicaps that undermined performance in similar circumstances, and ($iv$) may have misconfigured and underreported its baselines. We spell these out in Sections 3.1-3.4 — confirming even a fraction of specific concerns would put the top-line claims and conclusions of [1] in serious doubt.
Figure 1: Layouts from [6, Figure 2] with macros in red and standard cells in green, locations produced by RL (left) and RePlAce (right) for the ibm10 benchmark from [23]. Limiting macro locations to a coarse grid (left) leads to spreading of small macros (red squares on a grid) and elongates connecting wires: from 27.5 (right) to 44.1 (left) for ibm10 [6, Table 1]). High area utilization and many macros of different sizes make the ICCAD 2004 benchmarks [23] challenging compared to [1] and [2, page 43].
@@ 3.1 Unsubstantiated claims and
U7. For iterative optimization algorithms that improve results over time, fair comparisons show per testcase: ($i$) better quality metrics with equal runtime, ($ii$) better runtime with equal quality or ($iii$) wins for both. [1] offered no such evidence. In particular, if ML-based optimization is used with extraordinary compute resources, then so should be optimization by Simulated Annealing in its most competitive form.
@@ 4 Additional evidence
Months after the Nature publication [1], more data became available in [2, 4, 5, 6], followed by the first wave of controversial media coverage [54, 55, 56]. Nature editors released the peer review file for [1] including authors’ rebuttals: in the lengthy back-and-forth with reviewers [2] the authors assured reviewers that macro locations were not modified after placement by RL, confirming coarse-grid placement of macros. Among several contributions, [6] implemented the request of Nature Reviewer #3 [2] and benchmarked Google’s technique on 17 public chip-design examples [23]: prior methods decisively outperformed Google RL. Professors Patrick Madden (SUNY Binghamton) and Jens Lienig (TU Dresden) publicly expressed doubts about the Nature paper [54, 55]. As researchers noted gaps in the Google open-source release [4], such as the grouping (clustering) flow, Google engineers released more code (but not all), prompting more questions (see below). Another year passed, and [7, 8] expanded on the initial suspicions in several ways. They demonstrated that not limiting macro placement allows human designers and commercial EDA tools (separately) to outperform results produced by Google code [4]. [8, Table 2] estimated rank correlation of the proxy cost function optimized by RL to chip metrics used in [1, Table 1], and [8, Table 3] estimated the mean and standard deviation for chip metrics after RL-based optimization. We give a summary in Table 2, where rank correlations are low for all chip metrics, while TNS and WNS are noisy. Hence, the optimization of TNS and WNS in [1] relied on a flawed proxy and produced results in [1, Table 1] of dubious statistical significance. We note that $\sigma/|\mu|>0.5$ for TNS on Ariane-NG45 (also on BlackParrot-NG45 in [8, Table 3]). In the second round of critical media coverage, [1] was questioned by Profs. William Swartz (UT Dallas), Patrick Madden (SUNY Binghamton), and Moshe Vardi (Rice) [61, 64].
@@
Rank correlation to RL proxy cost 0.00 0.28 0.05 0.20 0.05
@@
Table 2: Evaluating the soundness of the proxy cost used with RL in [1] and the noisiness of reported chip metrics after RL-based optimization. We summarize data from [8, Table 2] on the Kendall rank correlation of chip metrics to the RL proxy cost and from [8, Tables 3 and 4] on statistics for chip metrics (only Ariane-NG45 design data are shown, but data for BlackParrot-NG45 shows similar trends). As expected, purely-additive metrics (standard-cell area, routed wirelength and chip power) exhibit low variance, but the TNS and WNS metrics, that measure timing-constraint violations, have high variance.
@@ 4.1 Methods
Undisclosed use of $(x,y)$ locations from commercial tools. Strong evidence and confirmation by Google engineers are mentioned in the UCSD paper [8] that [1] withheld a critical detail. When clustering the input netlist, CT merge in [4] read in a placement to restructure clusters based on locations. To produce $(x,y)$ locations of macros, [1] used initial $(x,y)$ locations of all circuit components (including macros!) produced by commercial EDA tools from Synopsys [4, Issue #25]. The lead authors of [1] confirmed using this step, claiming it was unimportant [60]. But it improved key metrics by 7-10% in [8]. So, the results in [1] needed algorithmic steps absent in [1], e.g., $(x,y)$ data from commercial software.
@@
Data leakage between training and test data? Per [1], “as we expose the policy network to a greater variety of chip designs, it becomes less prone to overfitting.” But Google Team 1 showed later in [5, Figure 7] that pre-training on “diverse TPU blocks” did not improve quality of results. Pre-training on “previous netlist versions” improved quality somewhat. Pre-training RL and evaluating it on similar designs could be a serious flaw in methodology of [1]. [^6] Such a methodology could help chip designers iterate on design changes faster, but that was not described in [1]. As Google did not release TPU designs or per-design statistics, we cannot compare training and test data.
@@ 4.3 Likely imitations
[1] did not disclose major limitations of its methods but promised success in broader combinatorial optimization. The Ariane design image in [1, Extended Data Figure 4] shows macro blocks of identical sizes: a potential limitation. Yet, [1] does not report basic statistics per TPU block: the number of macros and their shapes, design area utilization, and the fraction of area taken by macros. Based on [2, page 43] and the guidance from Google engineers to the authors of [8], it appears TPU blocks had area utilization on the order of 60%. Commercial chips are often denser, and can use macros of different sizes. Poor performance of Google RL on challenging public benchmarks from [23] used in [6, 8] (illustrated in Figure 1) suggests undisclosed limitations. Another possible limitation is poor handling of preplaced (fixed) macros, common in industry layouts, but not discussed in [1]. By interfering with preplaced macros, gridding (see H4) can impact usability in practice. Poor performance on public benchmarks may also be due to overfitting to proprietary TPU designs.
$\downarrow$ Designs / Tools $\rightarrow$ Google CT/RL Cadence CMP UCSD SA
@@ 5 Did [1] improve SOTA?
The Nature editorial [3] discussing [1] speculated that “this is an important achievement and will be a huge help in speeding up the supply chain.” But today, after evaluations and reproduction attempts at multiple chip-design and EDA companies, it is safe to conclude that no important achievement occurred because prior chip-design software, particularly from Cadence, produced better layouts faster [7, 8]. If this were known to the reviewers of [1] or to the public, the claims of improving TPU designs in [1] would be nonsensical. [1] claimed that humans produced better results than commercial EDA tools, but gave no substantiation. When license terms complicate publishing comparisons to commercial EDA tools, [^7] The lawsuit [59, 66] alleges that Google obtained better results with commercial tools before Nature submission. one compares to academic software and to other prior methods, with the proviso that small improvements are not compelling.
@@ 5.1 Reproduction attempts
  • Google Team 2 had access to TPU design blocks and evaluated the impact of pre-training in [6]. The impact was small at best. [^8] A consistent conclusion was reported in [5, Figure 7] by Google Team 1 — training on diverse designs does not help, and improvements are seen only when training on earlier versions of the same design. In December 2022 Dr. Jeff Dean, a then Google SVP and the most senior author of [1], also confirmed that RL did well without pre-training [58].
@@
As explained in [6], the implementation of Simulated Annealing used in [1] was handicapped. Removing the handicaps (in the same source code base) improved results. When properly implemented, SA produces better solutions than Google CT/RL [4] using less runtime, when both are given the same proxy cost function. This is shown consistently in [6, 8] on 17 widely used ICCAD 2004 benchmarks [23] and in [8] on several modern design benchmarks. Compared to Google CT/RL [4], SA consistently improves wirelength and power metrics. For circuit timing metrics TNS and WNS, SA produces less noisy results but comparable to RL’s results [8]. Recall that the proxy function optimized by SA and RL does not include timing metrics [1], making any claims of improvement in these metrics due to SA or RL dubious.